OpenStudy (anonymous):

Find the solution to the following system of equations. 2x + y = 7 x – y = -4

OpenStudy (anonymous):

here we go again... you can add both equation and eliminate y then solve for x...

OpenStudy (anonymous):

2x + y + x - y = 7 + 4 3x = 11 or x = 11/3

OpenStudy (anonymous):

you can use either of the two given equations to solve for y by plug-in x...

OpenStudy (anonymous):

easier to use 2nd equation: x - y = 4 11/3 - y = 4 or y = 11/3 - 4 y = (11 - 4(3))/3 y = (11 - 12)/3 y = -1/3

OpenStudy (anonymous):

you can check by plug-in both to equation 1 if both x and y satisfy the equation...

OpenStudy (anonymous):

2x + y = 7 2(11/3) + (-1/3) ?= 7 22/3 - 1/3 ?= 7 (22 - 1)/3 ?= 7 21/3 ?= 7 7 = 7.... check they are equal